OpenStudy (anonymous):

what would the tempurature be outside a jet flying at 40,000 ft? with a slope of -30 degrees/10,000ft and a y intercept of 40 degrees?

OpenStudy (dumbcow):

so the linear function of temperature and altitude is: T = 40 - 30h, where h is in 10,000 ft plug in h = 4 T = 40-30(4) = 40 - 120 = -80